OpenStudy (anonymous):

Pls help pretty Pls! will Medal! which percent is equivalent to 5 1/20? A. 5.05% B. 5.12% C. 505% D. 520%

OpenStudy (nikki_skellington):

I will help

OpenStudy (anonymous):

A

OpenStudy (anonymous):

medal?

OpenStudy (nikki_skellington):

no c

OpenStudy (anonymous):

Hm okay nikki so hows it c? and liam nice name and hows it a?

OpenStudy (nikki_skellington):

For sure it is C

OpenStudy (anonymous):

can u explain a little

OpenStudy (anonymous):

okay i'll just go with A i guess :(

OpenStudy (ilovemath14):

5 1/20=5.05 Percents cant have a decimal so the percent would be 5005% Which is C

OpenStudy (anonymous):

okay thanks than ill go with c! <3 <3

OpenStudy (ilovemath14):

Your welcome

OpenStudy (anonymous):

First write 1/20 and multiply both sides by 5, you will get 5/100 = 0.05. Add 5 to that. So the equation is 5 + 0.05 = ?